# Imprecise Markov Semigroups and their Ergodicity

### v4

@@ 5. Conclusion
In this work, we introduced the concept of an imprecise Markov semigroup, and studied its $\tilde{f}-$ergodic behavior, for some $\tilde{f}\in B(E)$, both when the state space $E$ is a Euclidean space or a Riemannian manifold, and when $E$ is allowed to be an arbitrary good measurable space.
In the future, we plan to forego the assumption that $\mathbf{Q}$ is compact, and work directly with nonlinear Markov semigroups (Kühn, 2021; Feng and Zhao, 2021). This will require a considerable effort, since a new theory involving the Bakry-Émery curvature condition associated with nonlinear Markov semigroups will need to be developed.
Before this (ambitious) goal, we plan to apply our findings to machine learning and computer vision, as we discussed in the Introduction, and to achieve robustness in reinforcement learning, where Markov processes are used for planning in autonomous agents. We also aim to extend to the imprecise case the types of inequalities put forth in Bakry et al. (2014, Chapters 6-9). In particular, we are interested in deriving imprecise optimal transport inequalities, similarly to (Caprio, 2024; Lorenzini et al., 2024), with the intent of applying them to distribution shift problems in machine learning and computer vision (Lin et al., 2023).
@@ Acknowledgements
We would like to acknowledge partial funding by the Army Research Office (ARO MURI W911NF2010080). We wish to express our gratitude to Mira Gordin for insightful discussions, particularly on how the gradient bound for a diffusion operator relates to Poincaré inequalities, and to Simone Cerreia-Vioglio and Paolo Perrone for their help in finding the suitable topology compatible with the partial order $\preceq^{\text{part}}_{\tilde{f}}$, and in shaping up the discussion after Lemma 6. We are also extremely grateful to Enrique Miranda for his suggestion to look into Krak et al. (2017, Example 6.2), to Andrea Aveni for his comments on an initial version of this work, and to Malena Español for her invitation to the Institute for Advanced Study, where the idea of the present paper germinated.
@@ Appendix A Proof of Theorem 10
We begin by pointing out that, by Lemma 6 and Definition 9, $\underline{\mathbf{P}}$ and $\overline{\mathbf{P}}$ are (precise) Markov semigroups, and so the operators $\underline{\Gamma}$, $\overline{\Gamma}$, $\underline{L}$ and $\overline{L}$ are well-defined.
We first focus on the diffusion Markov triple $(E,\overline{\mu},\overline{\Gamma})$. By Bakry et al. (2014, Proposition 3.3.11) and the fact that $\overline{\mathbf{P}}=(\overline{P}_{t})_{t\geq 0}$ is a (precise) Markov semigroup, we have that $\overline{L}$ is essentially self-adjoint. Pick any $f\in\mathcal{A}_{0}^{\overline{\mu}}$. The fact that $\overline{L}$ satisfies ESA, together with our assumption that it satisfies the curvature condition $CD(\rho_{\overline{L}},\infty)$, for some $\rho_{\overline{L}}\in\mathbb{R}$, implies – by Bakry et al. (2014, Corollary 3.3.19) – that for all $t\geq 0$, the following gradient bound holds
@@
Following van Handel (2016, Problem 2.7), taking the limit as $t\rightarrow\infty$ on both sides of (12) yields the following Poincaré inequality
@@
In addition, we know that
@@
Combining (14) with (13), we have that if function $f$ is such that $\overline{L}f=0$, then $\overline{\mathcal{E}}(f)=0$. In turn, this implies that $\text{Var}_{\overline{\mu}}(f)=0$, and so the function $f$ must be constant $\overline{\mu}$-a.e. Notice now that, since $\overline{L}$ satisfies the ESA property, $\mathcal{A}_{0}^{\overline{\mu}}$ is dense in $\mathcal{D}(\overline{\mathcal{E}})$ in the topology induced by the Dirichlet norm $\|\cdot\|_{\overline{\mathcal{E}}}$. [^18] This norm is recalled in Definition 18.(DMT4) in B.2. Thanks to this, (11)-(14) hold for all $f\in\mathcal{D}(\overline{\mathcal{E}})$ by an approximation argument. This implies that every $f\in\mathcal{D}(\overline{\mathcal{E}})$ such that $\overline{\Gamma}(f)=0$ is constant, and so that $\overline{L}$ is ergodic according to Bakry et al. (2014, Definition 3.1.11). Then, since $\overline{\mu}$ is finite, by Bakry et al. (2014, Proposition 3.1.13) we have that
@@
Appendix B Background Notions
@@
B.1. Background on Markov Semigroups
@@
$\mathcal{A}\ni\Gamma_{2}(f,g)\coloneqq\frac{1}{2}\left[L\Gamma(fg)-\Gamma(f,Lg)-\Gamma(Lf,g)\right],$ (18)
@@
As for the carré du champ operator $\Gamma$, we write $\Gamma_{2}(f,f)\equiv\Gamma_{2}(f)$ to lighten the notation. Unlike $\Gamma$, $\Gamma_{2}$ is not always positive. As we will see more formally in B.1.1, $\Gamma_{2}$ is positive if and only if the state space $E$ has “positive curvature”. This high-level intuition, together with the fact that many differential inequalities can be expressed via $\Gamma_{2}$ and $\Gamma$, justifies the following definition (Bakry et al., 2014, Definition 1.16.1).
@@ B.1.1. Riemannian Manifold
B.2. Background on Diffusion Markov Triples
@@
    The semigroup $\mathbf{P}=(P_{t})_{t\geq 0}$ is the symmetric semigroup with infinitesimal generator $L$ defined on its domain $\mathcal{D}(L)$. In general, it is sub-Markov, that is, $P_{t}(\mathbbm{1})\leq\mathbbm{1}$, for all $t\geq 0$, where $\mathbbm{1}$ is the constant function equal to $1$. For it to be Markov, $P_{t}(\mathbbm{1})=\mathbbm{1}$ must hold for all $t\geq 0$. Moreover, $\mathcal{A}_{0}$ is not necessarily dense in the domain $\mathcal{D}(L)$ with respect to the domain norm $\|f\|_{\mathcal{D}(L)}=[\|f\|_{2}^{2}+\|Lf\|_{2}^{2}]^{1/2}$. By (DMT8), we have that $\mu$ is invariant for $\mathbf{P}$. In addition, from the symmetry of $\Gamma$ we have that $\mu$ is reversible for $\mathbf{P}$.
@@
As we can see, the ESA property depends on the choice of $\mathcal{A}_{0}$. In addition, ESA holds for all elliptic operators on complete Riemannian manifolds where $\mathcal{A}_{0}$ is the class of smooth compactly supported functions. That is, it is always satisfied when $E$ is a Euclidean space or a complete Riemannian manifold. The other implication, though, need not hold – that is, there are ESA operators that are not elliptic as in B.1.1. As we shall see, ESA alone is enough to guarantee ergodicity à la Corollary 8.1.
@@
When $E$ is a Euclidean space or a complete Riemannian manifold, we have that $\mathcal{A}_{0}$ is the class of smooth ($\mathcal{C}^{\infty}$) compactly supported functionals on $E$, while $\mathcal{A}$ is the class of smooth ($\mathcal{C}^{\infty}$) functionals. We now present two concepts, connexity and weak hypo-ellipticity, that allow to “import” some of the geometric structure of the Riemannian manifolds considered in B.1.1 to the more general diffusion Markov triple setting (Bakry et al., 2014, Section 3.3.3). In addition, as we can see in the proof of Theorem 10, in the context of imprecise diffusion Markov tuples they are sufficient for the ESA property to hold.

### v5

@@ 6 Conclusion
In this work, we introduced the concept of an imprecise Markov semigroup $\mathbf{Q}$, and studied its ergodic behavior (i) when the state space $E$ is a Euclidean space or a Riemannian manifold and $\mathbf{Q}$ is compact; (ii) when $E$ is allowed to be an arbitrary good measurable space and $\mathbf{Q}$ is compact; and (iii) when $E$ is a generic Polish metric space with finite diameter, and $\mathbf{Q}$ is not assumed compact.
In future work, we plan to extend the results of Section 5 beyond the function class $\text{Lip}_{b}(E)$ and to relax the finite-diameter assumption (11). We also plan to extend our theory to explicitly model ambiguity in the initial distribution of a Markov process.
Furthermore, we aim to apply these results to machine learning and computer vision, thereby making the high-level discussion in the Introduction more concrete, and to develop robust reinforcement learning methods, where Markov processes are used for planning by autonomous agents.
We are also interested in extending to the imprecise setting the types of inequalities developed in Bakry et al. (2014, Chapters 6-9). In particular, we are interested in deriving imprecise optimal transport inequalities, in the spirit of (Caprio, 2024; Lorenzini et al., 2025), for application to distribution shift problems in machine learning and computer vision (Lin et al., 2023).
@@ Acknowledgements
We would like to acknowledge partial funding by the Army Research Office, grant ARO MURI W911NF2010080. We wish to express our gratitude to Mira Gordin for insightful discussions, particularly on how the gradient bound for a diffusion operator relates to Poincaré inequalities, and to Simone Cerreia-Vioglio and Paolo Perrone for their help in finding the suitable topology compatible with the partial order $\preceq^{\text{part}}_{\tilde{f}}$, and in shaping up the discussion after Lemma 7. We are also extremely grateful to Enrique Miranda for his suggestion to look into Krak et al. (2017, Example 6.2), to Andrea Aveni for his comments on an initial version of this work, and to two anonymous reviewers for their invaluable remark that greatly improved our manuscript. We are also indebted to Malena Español for her invitation to the Institute for Advanced Study, where the idea of the present paper germinated.
@@ Appendix A Proof of Theorem 12
We begin by pointing out that, as a consequence of Lemma 7 and Definition 11, $\underline{\mathbf{P}}$ and $\overline{\mathbf{P}}$ are (precise) Markov semigroups, and so the operators $\underline{\Gamma}$, $\overline{\Gamma}$, $\underline{L}$ and $\overline{L}$ are well-defined.
We first focus on the diffusion Markov triple $(E,\overline{\mu},\overline{\Gamma})$. By Bakry et al. (2014, Proposition 3.3.11) and $\overline{\mathbf{P}}=(\overline{P}_{t})_{t\geq 0}$ being a (precise) Markov semigroup, the fact that $\overline{L}$ is essentially self-adjoint follows from the connexity, weak hypo-ellipticity, and completeness assumptions. Pick any $f\in\mathcal{A}_{0}^{\overline{\mu}}$. The fact that $\overline{L}$ satisfies ESA, together with our assumption that it satisfies the curvature condition $CD(\rho_{\overline{L}},\infty)$, for some $\rho_{\overline{L}}\in\mathbb{R}$, implies – by Bakry et al. (2014, Corollary 3.3.19) – that for all $t\geq 0$, the following gradient bound holds
@@
Here the factor needs to be understood as $2t$ when $\rho_{\overline{L}}=0$. Following van Handel (2016, Theorem 2.35), taking the limit as $t\rightarrow\infty$ on both sides of (32) yields the following Poincaré inequality
@@
As in the proof of Theorem 10, here it is implicitly assumed that $\rho_{\overline{L}}>0$. In addition, we know that
@@
Combining (34) with (33), we have that if function $f$ is such that $\overline{L}f=0$, then $\overline{\mathcal{E}}(f)=0$. In turn, this implies that $\text{Var}_{\overline{\mu}}(f)=0$, and so the function $f$ must be constant $\overline{\mu}$-a.e. Notice now that, since $\overline{L}$ satisfies the ESA property, $\mathcal{A}_{0}^{\overline{\mu}}$ is dense in $\mathcal{D}(\overline{\mathcal{E}})$ in the topology induced by the Dirichlet norm $\|\cdot\|_{\overline{\mathcal{E}}}$. [^16] This norm is recalled in Definition 33.(DMT4) in C.2. Thanks to this, (31)-(34) hold for all $f\in\mathcal{D}(\overline{\mathcal{E}})$ by an approximation argument. This implies that every $f\in\mathcal{D}(\overline{\mathcal{E}})$ such that $\overline{\Gamma}(f)=0$ is constant, and so that $\overline{L}$ is ergodic according to Bakry et al. (2014, Definition 3.1.11). Then, since $\overline{\mu}$ is finite, by Bakry et al. (2014, Proposition 3.1.13) we have that
@@
Appendix B Measurability of Envelopes and Switching Operators
Under Assumption 1, for each $t\geq 0$ and bounded Borel $f$, the map $(x,\theta)\mapsto P_{t}^{\theta}f(x)$ is Borel measurable (by a standard monotone class argument, starting from indicators $\mathbbm{1}_{A}$). Now, let us denote by $\mathcal{F}^{\star}$ the universal completion of $\mathcal{F}$, and by $B^{\star}(E)$ the space of bounded $\mathcal{F}^{\star}$-measurable real-valued functionals on $E$. Since $f$ is $\mathcal{F}^{\star}$-measurable, it is measurable with respect to the completion of $p_{t}^{\theta}(x,\cdot)$ for every $(t,x,\theta)$; Hence we may extend $P_{t}^{\theta}$ to $B^{\star}(E)$ by
$P_{t}^{\theta}f(x)\coloneqq\int_{E}f(y)p_{t}^{\theta}(x,\mathrm{d}y),\qquad f\in B^{\star}(E),$
which is well-defined and $\mathcal{F}^{\star}$-measurable in $x$.
Consequently, for each fixed $t\geq 0$ and bounded Borel $f$, the envelope $x\mapsto\sup_{\theta\in\Theta}P_{t}^{\theta}f(x)$ is upper semianalytic, hence $\mathcal{F}^{\star}$-measurable. Likewise, the time-consistent switching operator introduced in Definition 20 is obtained by composing the one-step envelopes $\overline{T}$ along a finite partition of $[0,t]$ and then taking a supremum over partitions. [^17] Similarly for the lower Nisio semigroup in Definition 24. Since upper semianalytic functions are stable under application of Borel kernels – i.e. $g\mapsto(x\mapsto\int g(y)p(x,\text{d}y))$ – and under pointwise suprema, for bounded Borel $f$ each $T_{t}f$ is upper semianalytic and therefore $\mathcal{F}^{\star}$-measurable. For general $f\in B^{\star}(E)$ the same conclusion holds at the level of universal measurability, so all operators are interpreted as acting on $B^{\star}(E)$.
Appendix C Background Notions
@@
C.1 Background on Markov Semigroups
@@
$\mathcal{A}\ni\Gamma_{2}(f,g)\coloneqq\frac{1}{2}\left[L\Gamma(f,g)-\Gamma(f,Lg)-\Gamma(Lf,g)\right],$ (38)
@@
As for the carré du champ operator $\Gamma$, we write $\Gamma_{2}(f,f)\equiv\Gamma_{2}(f)$ to lighten the notation. Unlike $\Gamma$, $\Gamma_{2}$ is not always positive. As we will see more formally in C.1.1, $\Gamma_{2}$ is positive if and only if the state space $E$ has “positive curvature”. This high-level intuition, together with the fact that many differential inequalities can be expressed via $\Gamma_{2}$ and $\Gamma$, justifies the following definition (Bakry et al., 2014, Definition 1.16.1).
@@ C.1.1 Riemannian Manifold
C.2 Background on Diffusion Markov Triples
@@
    The semigroup $\mathbf{P}=(P_{t})_{t\geq 0}$ is the symmetric semigroup with infinitesimal generator $L$ defined on its domain $\mathcal{D}(L)$. In general, it is sub-Markov, that is, $P_{t}(\mathbbm{1})\leq\mathbbm{1}$, for all $t\geq 0$, where $\mathbbm{1}$ is the constant function equal to $1$. For it to be Markov, $P_{t}(\mathbbm{1})=\mathbbm{1}$ must hold for all $t\geq 0$. Moreover, $\mathcal{A}_{0}$ is not necessarily dense in the domain $\mathcal{D}(L)$ with respect to the domain norm $\|f\|_{\mathcal{D}(L)}=[\|f\|_{2}^{2}+\|Lf\|_{2}^{2}]^{1/2}\equiv\|f\|_{\mathbb{L}^{2}(\mu)}+\|Lf\|_{\mathbb{L}^{2}(\mu)}$. By (DMT8), we have that $\mu$ is invariant for $\mathbf{P}$. In addition, from the symmetry of $\Gamma$ we have that $\mu$ is reversible for $\mathbf{P}$.
@@
Before recalling adjoints and essential self-adjointness, we briefly fix notation on the graph norm and closedness for unbounded operators, since we will use approximation on $\mathcal{D}(L)$ through a core.
Let $(H,\|\cdot\|_{H})$ be a Banach space and let $A:\mathcal{D}(A)\subset H\to H$ be a linear operator. The graph of $A$ is
$\mathrm{Graph}(A)\coloneqq\{(f,Af):f\in\mathcal{D}(A)\}\subset H\times H,$
and the associated graph norm on $\mathcal{D}(A)$ is
$\|f\|_{\mathcal{D}(A)}\coloneqq\|f\|_{H}+\|Af\|_{H}.$
We recall that $A$ is closed if and only if $\mathrm{Graph}(A)$ is closed in $H\times H$. Equivalently, $A$ is closed if and only if $(\mathcal{D}(A),\|\cdot\|_{\mathcal{D}(A)})$ is a Banach space. Indeed, the map $f\mapsto(f,Af)$ is an isometric embedding of $\mathcal{D}(A)$ into $H\times H$ endowed with $\|(u,v)\|_{H\times H}\coloneqq\|u\|_{H}+\|v\|_{H}$, and completeness is equivalent to closedness of the range. In our setting, $H=\mathbb{L}^{2}(\mu)$ and the generator $L$ of the strongly continuous Markov semigroup $(P_{t})_{t\geq 0}$ is closed; Hence, $\mathcal{D}(L)$ is complete for the graph norm $\|\cdot\|_{\mathcal{D}(L)}$.
When we write that $\mathcal{A}_{0}$ is dense in $\mathcal{D}(L)$ for the graph norm, we mean that $\mathcal{A}_{0}$ is a core for $L$: For every $f\in\mathcal{D}(L)$ there exists $(f_{n})_{n}\subset\mathcal{A}_{0}$ such that
$\|f_{n}-f\|_{\mathbb{L}^{2}(\mu)}+\|Lf_{n}-Lf\|_{\mathbb{L}^{2}(\mu)}\rightarrow 0.$
This approximation is precisely what allows us to extend identities first proved on $\mathcal{A}_{0}$ (e.g. integration by parts and the Poincaré inequality) to all $f\in\mathcal{D}(L)$ by continuity.
@@
As we can see, the ESA property depends on the choice of $\mathcal{A}_{0}$. In addition, ESA holds for all elliptic operators on complete Riemannian manifolds where $\mathcal{A}_{0}$ is the class of smooth compactly supported functions. That is, it is always satisfied when $E$ is a Euclidean space or a complete Riemannian manifold. The other implication, though, need not hold – that is, there are ESA operators that are not elliptic as in C.1.1. As we shall see, ESA alone is enough to guarantee ergodicity à la Corollary 10.1.
@@
When $E$ is a Euclidean space or a complete Riemannian manifold, we have that $\mathcal{A}_{0}$ is the class of smooth ($\mathcal{C}^{\infty}$) compactly supported functionals on $E$, while $\mathcal{A}$ is the class of smooth ($\mathcal{C}^{\infty}$) functionals. We now present two concepts, connexity and weak hypo-ellipticity, that allow to “import” some of the geometric structure of the Riemannian manifolds considered in C.1.1 to the more general diffusion Markov triple setting (Bakry et al., 2014, Section 3.3.3). In addition, as we can see in the proof of Theorem 12, in the context of imprecise diffusion Markov tuples they are sufficient for the ESA property to hold.